OpenStudy (anonymous):

Simplify: 3u + 7v + 4u – 2v this is so confusing ;--;? I am extremely bad in math.

HanAkoSolo (jamierox4ev3r):

Hello there! In problems like these, we want to look for like terms. In this problem, 3u and 4u are like terms. 7v and -2v are also like terms. Following me so far?

OpenStudy (anonymous):

yes @Jamierox4ev3r

HanAkoSolo (jamierox4ev3r):

Excellent. To solve this equation, you're going to want to add the like terms together. For example, 3u+4u=7u. Can you do the other set of like terms for me? @izy_valle

OpenStudy (anonymous):

7v + 2v = 9v ? @Jamierox4ev3r

HanAkoSolo (jamierox4ev3r):

Close! However, you are not adding 7v and 2v, but rather 7v + (-2v). This can be rewritten as 7v-2v. Can you tell me now?

OpenStudy (anonymous):

5v? :)

HanAkoSolo (jamierox4ev3r):

Yes! So your final answer is..?

HanAkoSolo (jamierox4ev3r):

In the future, think about it this way: \(\large \color{green}{3u} \color{blue}{+ 7v }~ \color{green}{+4u} ~\color{blue}{– 2v}\) \(\large \color{green}{3u+4u=7u}\) \(\large \color{blue}{7v-2v=5v}\) Final answer: \(\large \color{green}{7u}+~\color{blue}{5v}\)

HanAkoSolo (jamierox4ev3r):

The greens are added together. Once you have that, you subtract 2v from 7v, since 2v has a subtraction sign in front of it. Once you have these terms, you add them together and get your final answer. Does this make sense?

HanAkoSolo (jamierox4ev3r):

@izy_valle you there?